One side of a rectangular stage is 2 meters longer than the other. If the diagonal is 10 meters, then what are the lengths of the sides?
Answer by checkley77(12844)

The 2 sides and the diagonal form a right triangle. Thus we have the formula: a^2+b^2=c^2
x^2+(x+2)^2=10^2
x^2+x^2+4x+4=100
2x^2+4x+4-100=0
2x^2+4x-96=0
2(x^2+2x-48)=0
2(x+8)(x-6)=0
x-6=0
x=6 answer for the shorter side.
6+2=8 for the longer side.
Proof:
6^2+8^2=10^2
36=64=100
100=100